OpenStudy (anonymous):

A rectangle is shown. The length is eight meters and the width is one meter. whats the area

OpenStudy (angela210793):

Al*w=8*1=8

OpenStudy (angela210793):

8m^2

OpenStudy (anonymous):

area=l*w

OpenStudy (amistre64):

\[\int_{0}^{1}\int_{0}^{8}dl\ dw\]

OpenStudy (akshay_budhkar):

lol amistre!!! Scaring a young child han?

OpenStudy (amistre64):

its a hobby :)

OpenStudy (akshay_budhkar):

What a hobby lol!! u do it with ur children too(just asking)